OpenStudy (deercult):

Solve the exponential equation. 125^(8x-2) = 150

OpenStudy (misty1212):

HI (again)

OpenStudy (deercult):

Hullo again! : P

OpenStudy (misty1212):

change of base formula for this one \[\huge b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (misty1212):

so \[125^{8x-2}=150\\ 8x-2=\frac{\ln(150)}{\ln(125)}\] then algebra

OpenStudy (deercult):

0.3797. It's an available answer! :>